# Towards Explainable Spoofed Speech Attribution and Detection: a Probabilistic Approach for Characterizing Speech Synthesizer Components

### v2

@@ Abstract
We propose an explainable probabilistic framework for characterizing spoofed speech by decomposing it into probabilistic attribute embeddings. Unlike raw high-dimensional countermeasure embeddings, which lack interpretability, the proposed probabilistic attribute embeddings aim to detect specific speech synthesizer components, represented through high-level attributes and their corresponding values. We use these probabilistic embeddings with four classifier back-ends to address two downstream tasks: spoofing detection and spoofing attack attribution. The former is the well-known bonafide-spoof detection task, whereas the latter seeks to identify the source method (generator) of a spoofed utterance. We additionally use Shapley values, a widely used technique in machine learning, to quantify the relative contribution of each attribute value to the decision-making process in each task. Results on the ASVspoof2019 dataset demonstrate the substantial role of duration and conversion modeling in spoofing detection; and waveform generation and speaker modeling in spoofing attack attribution. In the detection task, the probabilistic attribute embeddings achieve $99.7\%$ balanced accuracy and $0.22\%$ equal error rate (EER), closely matching the performance of raw embeddings ($99.9\%$ balanced accuracy and $0.22\%$ EER). Similarly, in the attribution task, our embeddings achieve $90.23\%$ balanced accuracy and $2.07\%$ EER, compared to $90.16\%$ and $2.11\%$ with raw embeddings. These results demonstrate that the proposed framework is both inherently explainable by design and capable of achieving performance comparable to raw CM embeddings.
@@ 1 Introduction
Advances in deep generative modeling [1] have paved way to flexible generation of high-quality synthetic media. Despite the many useful applications, this ‘deepfake[2] technology has raised concerns in its exploitation, whether for misinformation spreading in social media, defaming a targeted person, spoofing biometric systems [3] or conducting advanced social engineering attacks for financial gains, to name a few. Synthetic media targeted specifically against biometric recognition are forms of presentation attacks (spoofing attacks) [4]. In this study, our focus is on speech. The two well-known technologies of speech synthesis (text-to-speech, TTS) [5] and voice conversion (VC) [6] can be used for controlled generation and speaker identity conversion of speech utterances, respectively. While the past-generation TTS and VC technology was tailored for small and closed-world speaker populations, recent zero-shot [7] approaches make it possible, in principle, to reproduce speech in anyone’s voice based on a small amount of training data (required at the generation time only). Lately, a number of alerting social engineering attacks that involve the use of these technologies have been reported [8, 9, 10].
@@
When we know that $X$ actually originates from one of the attacks in $\mathcal{A}$, the spoofing attack attribution is said to have a closed set of spoofing attacks; otherwise an open set. Our study is not the first to address attack attribution, either in general multimedia [24, 25] or in speech (related studies are reviewed in Section 2). Existing approaches in the literature are primarily focused on predictive performance rather than interpretability. In contrast, we propose a novel probabilistic attribution framework (detailed in Section 3) which is designed to be inherently explainable. An initial study this method has been reported in our recent work [26], where the study was limited to using a classic decision tree classifier [27]. While decision trees provided a degree of interpretability, their predictive performance can be constrained. Moreover, the earlier work addressed the spoofing attack attribution task using only two attack. The present study substantially expands upon  [26] both at conceptual, theoretical framework and experimental levels. We incorporate a diverse range of interpretable back-end classifiers, including naive Bayes [28] (specifically tailored to handle probabilistic attribute features), logistic regression [29], and support vector machines (SVM) [30]. These classifiers bring complementary strengths, such as probabilistic reasoning (naive Bayes), robust classification (logistic regression), and the ability to handle decision boundaries with maximum margin for improved generalizability (SVM). This diversification enhances both the interpretability and predictive performance of the framework. Furthermore, to achieve better generalizability, this study introduces a new custom protocol on the ASVspoof2019 dataset, expanding the scope beyond the earlier focus on two known attack types to include all $17$ types of attacks for spoofing attack attribution.
@@
Figure 2: A tentative structure of spoofed speech generation, VC: voice conversion, E2E: End to End, TTS: text to speech.
@@
In this study, our intention is to demonstrate how different spoofing attack algorithms can be broken down into their elementary building blocks, represented through a high-level probabilistic feature description. In specific, our aim is to characterize spoofing attacks through abstract discrete probabilistic attributes. As for the ‘discrete’, random events that can be named, enumerated, and explained to others are convenient to us as humans. Much of the basic sciences (and everyday communication) seeks to explain complicated concepts through simplifications. As for the ‘probabilistic’, real-world physical objects, including acoustic speech waveforms, are rarely ‘either-this-or-that’ due to inaccuracies in measuring or observing the relevant cues; probabilities provide a natural way to quantify uncertainty in the features, allowing one to postpone the final classification decision to a later stage. Finally, ‘attribute’ is essentially a synonym for ‘feature’, but intended to emphasize the notion of the features to bear clear semantics to humans.
@@
Apart from the subtle technical modeling choice on attribute format, the more critical difference underlines in the task (speaker characterization in [33, 34, 35]; deepfake characterization in the present work). Speaker characterization leverages audio data that originates from a biological system—the human vocal production system driven by neural and muscular activities that consists of the lungs, the larynx, and the supralaryngeal vocal tract system. As such, motivation and design of attributes for speaker characterization can benefit from the established speech science that covers domains such as physiology, biomechanics, aerodynamics, acoustics, digital speech modeling, phonetics and phonology. In contrast, deepfakes originate from artificial systems (computer programs that implement specific synthesis or conversion models). What are the relevant attributes of such systems — how should we characterize spoofed and deepfake speech? Whereas the field of speech synthesis has a long history and its established system categorizations (e.g. concatenative, statistical, and neural TTS), there has been little focus towards the attribution task. In short, we require thorough re-thinking of what the explainable attributes for categorizing, classifying and explaining different spoofing attack algorithms should be. With deepfakes being still relatively recently security threat, and with major focus of defense solutions having been on the binary ’real-fake’ discrimination task, the work in more refined categorization of spoofed speech generators (including their detection) is currently taking only its infant steps. By no means the present work is intended to be a full description of spoofed speech characterization but, rather, a modest conceptual step forward. Besides forensic ‘speaker profiling’, an already established topic [37], we expect future forensic speech analysts to benefit from ‘deepfake profiling’ or ‘attack fingerprinting’ methodologies. To the best of the authors’ knowledge, however, no such standardized approaches are available yet.
@@
The second category attempts to explain spoofed speech by tracing its source of generation, commonly referred to as spoofing attack attribution or source tracing, as mentioned in Section 1. One of the earliest investigations in this direction [38] utilized bispectral magnitude and phase features to differentiate between spoofed and bonafide speech sources. Building on this, [39] explored long-term and short-term parameterization using linear prediction (LP) residuals and their periodicity to detect spoofing attributes, highlighting the importance of time-domain feature analysis. Open-set attribution was investigated in [40] through a one-class thresholding approach combined with a support vector machine (SVM). Similarly, [41] addressed open-set attribution using an ensemble semi-supervised multi-class strategy based on convolutional neural networks (CNNs). Further advances in spoofing attack attribution include approaches based on the popular Transformer models. For instance, [42] proposed a patchout spectrogram-based Transformer framework, while [43] introduced a compact attribution transformer with a poly-1 loss function [56] to enhance open-set performance. Extending this direction, [44] explored self-supervised Transformer models for spoofing attack attribution. Another study [45] examined low-level acoustic features such as fundamental frequency, distortion, shimmer, and speaker gender, combined with neural embeddings to improve attack attribution. Moreover, [53] demonstrated the role of prosodic signatures captured in speech pre-trained models, while [54] explored sub-band time-frequency features with Transformers for spoofing attack attribution. Recently, [46] employed Rawnet [57], Whisper [58], and self-supervised learning techniques [59] to trace spoofed speech sources, focusing on attributes such as input type, acoustic model, and vocoder [59]. Another approach in [47] utilized Rawnet and ResNet34 architectures with multi-task learning (joint attribute value classification and spoofing detection) to address source tracing alongside spoofing detection. Furthermore, [55] introduced neural codec source tracing and demonstrated its effectiveness in both open-set and closed-set scenarios, leveraging mel light CNNs (mel-LCNN) [60], audio anti-spoofing using integrated spectro-temporal graph attention network (AASIST) [61], and wav2vec-AASIST [62] architectures. The growing interest in this field is evident from the upcoming special session on Source tracing: The origins of synthetic or manipulated speech [^1] https://www.interspeech2025.org/special-sessions at Interspeech 2025, which aims to bring together researchers working on spoofing attribution techniques. Additionally, as summarized in Table 1, most prior works utilized publicly available datasets such as ASVspoof2019 [63], Defense Advanced Research Projects Agency Semantic Forensics (DARPA SemaFor) [64], TIMIT-TTS [65], Open Dataset of Synthetic Speech (ODSS) [66], Fake or Real (FoR) [67], Chinese fake audio detection (CFAD) [68] and Multi-Language Audio Anti-Spoof (MLAAD) [69] for evaluating the performance of their methods.
@@ 3.1 Attributes of spoofed speech generators
We aim to characterize spoofed speech using explainable probabilistic attribute embeddings, drawing inspiration from the processes involved in its generation. To this end, we decompose spoofed speech generation to its underlying attributes hierarchically, as illustrated in Figure 2. First, note that spoofed utterances are typically produced through voice conversion (VC) or text-to-speech (TTS) approaches, or their hybrids (TTS followed by VC) [47, 46, 63]. These methods are generally designed and implemented through one of the two main approaches. The first approach combines a number of clearly identified sub-components—such as an intermediate acoustic feature predictor and a vocoder—in a cascade. Since the individual sub-components (modules) and their functions are clearly understood and known at the training and inference (speech generation) stages, we refer to spoofed speech generators in this category modular. This is in contrast to the second approach based on end-to-end (E2E) [65] models that map either a text passage (in TTS) or a source speaker utterance (in VC) directly to a predicted waveform. The design philosophy of E2E approaches is to use a model optimized ‘on one go’, as opposed to constructing it from separately-optimized sub-components. Note that while E2E models also consist of cascaded components—namely, the layers in a deep neural network that pass intermediate latent feature representations to the next layer(s)—the key distinction here is whether the individual sub-components have a clearly-identified, explainable meaning to their function or not. A ’Wavenet vocoder’, for instance, might be argued to be more meaningful than ‘layer number 26’.
@@
The modules used in constructing a complete TTS or VC system themselves can further be characterized with a number of pre-defined high-level design choices, or attributes. The values of attributes determine the design choices used. For instance, the intermediate acoustic feature predictor operates on specific type of input data (text or speech) and produces particular type of outputs (type of predicted features) expected by the subsequent vocoder module. The intermediate acoustic feature predictor may use a specific duration model or a particular target speaker representation (such as one-hot vector or an external speaker embedding vector for zero-shot type of TTS). Assigning specific values for both the feature predictor and the vocoder attributes defines a particular TTS or VC system design.
@@
As a concrete example, consider one of the widely used modular TTS approaches described in  [71]. This approach consists of a so-called ‘Tacotron2’ acoustic feature predictor and a ‘WaveNet’ vocoder [72]. The former assumes character embedding inputs that are passed through layers of convolutional neural network, long-short term memory (LSTM), and bidirectional LSTM layers to predict a mel-spectrogram. A WaveNet vocoder conditioned on the mel-spectrogram is then used for generating the audio waveform. This approach can further be extended, as reported in  [73], to support multiple speakers by conditioning the mel-spectrogram generation by speaker embeddings. Hence, we might characterize the ‘Tacatron2+WaveNet’ TTS the following attribute-value pairs:
        input type          = text
        character encoder   = 1-hot
        feature prediction  = CNN+LSTM
        speaker embedding   = x-vector
        output features     = mel-spectrogram
        vocoder             = waveNet
@@ 3.3 Back-end classifiers
forms the basis for optimal decisions (decisions that minimize classification error probability). Here, $P(\bm{\mathrm{x}})\triangleq\sum_{j=1}^{N}\pi_{j}P_{j}(\bm{\mathrm{x}})$ is a normalizer that ensures that a valid posterior probability distribution results. Given the initially uncertain knowledge of the state $\mathcal{A}$, the Bayes’ rule provides a way to update the knowledge of the attack label. The prior $\pi_{i}$ and the likelihood $P_{i}(\bm{\mathrm{x}})$ indicate, respectively, the overall prevalence and typicality of an attack algorithm $\mathcal{A}_{i}$. For simplicity, we assume flat prior $\pi_{i}=1/C,\;\forall i=1,\dots,C$ and focus on modeling the likelihood terms.
@@
where $S_{\ell,m}\triangleq\sum_{n=1}^{N}\rho_{\ell,m}^{(n)}$ denotes the number of training observations with the value of 1 for attribute $m$ in attribute $\ell$, $N$ denoting the total number of utterances for a specific attack. The interested reader is pointed to the Appendix for the derivation, and to our public repository [^2] https://github.com/Manasi2001/Spoofed-Speech-Attribution for a reference implementation.

### v3

@@ Abstract
We propose an explainable probabilistic framework for characterizing spoofed speech by decomposing it into probabilistic attribute embeddings. Unlike raw high-dimensional countermeasure embeddings, which lack interpretability, the proposed probabilistic attribute embeddings aim to detect specific speech synthesizer components, represented through high-level attributes and their corresponding values. We use these probabilistic embeddings with four classifier back-ends to address two downstream tasks: spoofing detection and spoofing attack attribution. The former is the well-known bonafide-spoof detection task, whereas the latter seeks to identify the source method (generator) of a spoofed utterance. We additionally use Shapley values, a widely used technique in machine learning, to quantify the relative contribution of each attribute value to the decision-making process in each task. Results on the ASVspoof2019 dataset demonstrate the substantial role of waveform generator, conversion model outputs, and inputs in spoofing detection; and inputs, speaker, and duration modeling in spoofing attack attribution. In the detection task, the probabilistic attribute embeddings achieve $99.7\%$ balanced accuracy and $0.22\%$ equal error rate (EER), closely matching the performance of raw embeddings ($99.9\%$ balanced accuracy and $0.22\%$ EER). Similarly, in the attribution task, our embeddings achieve $90.23\%$ balanced accuracy and $2.07\%$ EER, compared to $90.16\%$ and $2.11\%$ with raw embeddings. These results demonstrate that the proposed framework is both inherently explainable by design and capable of achieving performance comparable to raw CM embeddings.
@@ 1 Introduction
Advances in deep generative modeling [1] have paved way to flexible generation of high-quality synthetic media. Despite the many useful applications, this ‘deepfake[2] technology has raised concerns in its exploitation, whether for misinformation spreading in social media, defaming a targeted person, spoofing biometric systems [3] or conducting advanced social engineering attacks for financial gains, to name a few. Synthetic media targeted specifically against biometric recognition are forms of presentation attacks (spoofing attacks) [4]. In this study, our focus is on speech. The two well-known technologies of speech synthesis (text-to-speech, TTS) [5] and voice conversion (VC) [6] can be used for controlled generation and speaker identity conversion of speech utterances, respectively. While the past-generation TTS and VC technology was tailored for small and closed-world speaker populations, recent zero-shot [7] approaches make it possible, in principle, to reproduce speech in anyone’s voice based on a small amount of training data (required at the generation time only). Lately, a number of alarming social engineering attacks that involve the use of these technologies have been reported [8, 9, 10].
@@
When we know that $X$ actually originates from one of the attacks in $\mathcal{A}$, the spoofing attack attribution is said to have a closed set of spoofing attacks; otherwise an open set. Our study is not the first to address attack attribution, either in general multimedia [24, 25] or in speech (related studies are reviewed in Section 2). Existing approaches in the literature are primarily focused on predictive performance rather than interpretability. In contrast, we propose a novel probabilistic attribution framework (detailed in Section 3) which is designed to be inherently explainable. An initial study this method has been reported in our recent work [26], where the study was limited to using a classic decision tree classifier [27]. While decision trees provided a degree of interpretability, their predictive performance can be constrained. Moreover, the earlier work addressed the spoofing attack attribution task using only two attacks. The present study substantially expands upon  [26] both at the conceptual, theoretical framework, and experimental levels. We incorporate a diverse range of interpretable back-end classifiers, including naive Bayes [28] (specifically tailored to handle probabilistic attribute features), logistic regression [29], and support vector machines (SVM) [30]. These classifiers bring complementary strengths, such as probabilistic reasoning (naive Bayes), robust classification (logistic regression), and the ability to handle decision boundaries with maximum margin for improved generalizability (SVM). This diversification enhances both the interpretability and predictive performance of the framework. Furthermore, to achieve better generalizability, this study introduces a new custom protocol on the ASVspoof2019 dataset, expanding the scope beyond the earlier focus on two known attack types to include all $17$ types of attacks for spoofing attack attribution.
@@
Refer to caption .
Figure 2: A tentative structure of spoofed speech generation, VC: voice conversion, E2E: End to End, TTS: text to speech, Waveform Gen.: Waveform Generator
@@
In this study, our intention is to demonstrate how different spoofing attack algorithms can be broken down into their elementary building blocks, represented through a high-level probabilistic feature description. Specifically, our aim is to characterize spoofing attacks through abstract discrete probabilistic attributes. As for the ‘discrete’, random events that can be named, enumerated, and explained to others are convenient to us as humans. Much of the basic sciences (and everyday communication) seeks to explain complicated concepts through simplifications. As for the ‘probabilistic’, real-world physical objects, including acoustic speech waveforms, are rarely ‘either-this-or-that’ due to inaccuracies in measuring or observing the relevant cues; probabilities provide a natural way to quantify uncertainty in the features, allowing one to postpone the final classification decision to a later stage. Finally, ‘attribute’ is essentially a synonym for ‘feature’, but intended to emphasize the notion of the features to bear clear semantics to humans.
@@
Apart from the subtle technical modeling choice on attribute format, the more critical difference is underlined in the task (speaker characterization in [33, 34, 35]; deepfake characterization in the present work). Speaker characterization leverages audio data that originates from a biological system—the human vocal production system driven by neural and muscular activities that consists of the lungs, the larynx, and the supralaryngeal vocal tract system. As such, motivation and design of attributes for speaker characterization can benefit from the established speech science that covers domains such as physiology, biomechanics, aerodynamics, acoustics, digital speech modeling, phonetics and phonology. In contrast, deepfakes originate from artificial systems (computer programs that implement specific synthesis or conversion models). What are the relevant attributes of such systems — how should we characterize spoofed and deepfake speech? Whereas the field of speech synthesis has a long history and its established system categorizations (e.g. concatenative, statistical, and neural TTS), there has been little focus towards the attribution task. In short, we require thorough re-thinking of what the explainable attributes for categorizing, classifying and explaining different spoofing attack algorithms should be. With deepfakes being still a relatively recent security threat, and with major focus of defense solutions having been on the binary ’real-fake’ discrimination task, the work in more refined categorization of spoofed speech generators (including their detection) is currently taking only its infant steps. By no means the present work is intended to be a full description of spoofed speech characterization but, rather, a modest conceptual step forward. Besides forensic ‘speaker profiling’, an already established topic [37], we expect future forensic speech analysts to benefit from ‘deepfake profiling’ or ‘attack fingerprinting’ methodologies. To the best of the authors’ knowledge, however, no such standardized approaches are available yet.
@@
The second category attempts to explain spoofed speech by tracing its source of generation, commonly referred to as spoofing attack attribution or source tracing, as mentioned in Section 1. One of the earliest investigations in this direction [38] utilized bispectral magnitude and phase features to differentiate between spoofed and bonafide speech sources. Building on this, [39] explored long-term and short-term parameterization using linear prediction (LP) residuals and their periodicity to detect spoofing attributes, highlighting the importance of time-domain feature analysis. Open-set attribution was investigated in [40] through a one-class thresholding approach combined with a support vector machine (SVM). Similarly, [41] addressed open-set attribution using an ensemble semi-supervised multi-class strategy based on convolutional neural networks (CNNs). Further advances in spoofing attack attribution include approaches based on the popular Transformer models. For instance, [42] proposed a patchout spectrogram-based Transformer framework, while [43] introduced a compact attribution transformer with a poly-1 loss function [56] to enhance open-set performance. Extending this direction, [44] explored self-supervised Transformer models for spoofing attack attribution. Another study [45] examined low-level acoustic features such as fundamental frequency, distortion, shimmer, and speaker gender, combined with neural embeddings to improve attack attribution. Moreover, [53] demonstrated the role of prosodic signatures captured in speech pre-trained models, while [54] explored sub-band time-frequency features with Transformers for spoofing attack attribution. Recently, [46] employed Rawnet [57], Whisper [58], and self-supervised learning techniques [59] to trace spoofed speech sources, focusing on attributes such as input type, acoustic model, and waveform generator [59]. Another approach in [47] utilized Rawnet and ResNet34 architectures with multi-task learning (joint attribute value classification and spoofing detection) to address source tracing alongside spoofing detection. Furthermore, [55] introduced neural codec source tracing and demonstrated its effectiveness in both open-set and closed-set scenarios, leveraging mel light CNNs (mel-LCNN) [60], audio anti-spoofing using integrated spectro-temporal graph attention network (AASIST) [61], and wav2vec-AASIST [62] architectures. The growing interest in this field is evident from the upcoming special session on Source tracing: The origins of synthetic or manipulated speech [^1] https://www.interspeech2025.org/special-sessions at Interspeech 2025, which aims to bring together researchers working on spoofing attribution techniques. Additionally, as summarized in Table 1, most prior works utilized publicly available datasets such as ASVspoof2019 [63], Defense Advanced Research Projects Agency Semantic Forensics (DARPA SemaFor) [64], TIMIT-TTS [65], Open Dataset of Synthetic Speech (ODSS) [66], Fake or Real (FoR) [67], Chinese fake audio detection (CFAD) [68] and Multi-Language Audio Anti-Spoof (MLAAD) [69] for evaluating the performance of their methods.
@@ 3.1 Attributes of spoofed speech generators
We aim to characterize spoofed speech using explainable probabilistic attribute embeddings, drawing inspiration from the processes involved in its generation. To this end, we decompose spoofed speech generation to its underlying attributes hierarchically, as illustrated in Figure 2. First, note that spoofed utterances are typically produced through voice conversion (VC) or text-to-speech (TTS) approaches, or their hybrids (TTS followed by VC) [47, 46, 63]. These methods are generally designed and implemented through one of the two main approaches. The first approach combines a number of clearly identified sub-components—such as an intermediate acoustic feature predictor and a waveform generator—in a cascade. Since the individual sub-components (modules) and their functions are clearly understood and known at the training and inference (speech generation) stages, we refer to spoofed speech generators in this category modular. This is in contrast to the second approach based on end-to-end (E2E) [65] models that map either a text passage (in TTS) or a source speaker utterance (in VC) directly to a predicted waveform. The design philosophy of E2E approaches is to use a model optimized ‘in one go’, as opposed to constructing it from separately-optimized sub-components. Note that while E2E models also consist of cascaded components—namely, the layers in a deep neural network that pass intermediate latent feature representations to the next layer(s)—the key distinction here is whether the individual sub-components have a clearly-identified, explainable meaning to their function or not. A ’WaveNet waveform generator’, for instance, might be argued to be more meaningful than ‘layer number 26’.
@@
The modules used in constructing a complete TTS or VC system themselves can further be characterized with a number of pre-defined high-level design choices, or attributes. The values of attributes determine the design choices used. For instance, the intermediate acoustic feature predictor operates on specific type of input data (text or speech) and produces particular type of outputs (type of predicted features) expected by the subsequent waveform generator module. The intermediate acoustic feature predictor may use a specific duration model or a particular target speaker representation (such as one-hot vector or an external speaker embedding vector for zero-shot type of TTS). Assigning specific values for both the feature predictor and the waveform generator attributes defines a particular TTS or VC system design.
@@
As a concrete example, consider one of the widely used modular TTS approaches described in  [71]. This approach consists of a so-called ‘Tacotron2’ acoustic feature predictor and a ‘WaveNet’ waveform generator [72]. The former assumes character embedding inputs that are passed through layers of convolutional neural network, long-short term memory (LSTM), and bidirectional LSTM layers to predict a mel-spectrogram. A WaveNet waveform generator conditioned on the mel-spectrogram is then used for generating the audio waveform. This approach can further be extended, as reported in  [73], to support multiple speakers by conditioning the mel-spectrogram generation by speaker embeddings. Hence, based on the structure shown in Figure 2, the ‘Tacotron2+WaveNet’ TTS system can be characterized using the following attribute-value pairs:
        input                                = text
        input processor(character encoder)   = 1-hot
        duration                             = None
        conversion                           = CNN+LSTM
        speaker                              = x-vector
        outputs                              = mel-spectrogram
        waveform generator                   = waveNet
@@ 3.3 Back-end classifiers
forms the basis for optimal decisions (decisions that minimize classification error probability). Here,
$P(\bm{\mathrm{x}})\triangleq\sum_{j=1}^{N}\pi_{j}P_{j}(\bm{\mathrm{x}}),$ (4)
is a normalizer that ensures that a valid posterior probability distribution results. Given the initially uncertain knowledge of the state $\mathcal{A}$, the Bayes’ rule provides a way to update the knowledge of the attack label. The prior $\pi_{i}$ and the likelihood $P_{i}(\bm{\mathrm{x}})$ indicate, respectively, the overall prevalence and typicality of an attack algorithm $\mathcal{A}_{i}$. For simplicity, we assume flat prior $\pi_{i}=1/C,\;\forall i=1,\dots,C$ and focus on modeling the likelihood terms.
@@
where $S_{\ell,m}\triangleq\sum_{n=1}^{N}\rho_{\ell,m}^{(n)}$ denotes the number of training observations with the value of 1 for attribute $m$ in attribute $\ell$, $N$ denoting the total number of utterances for a specific attack. The interested reader is pointed to Appendix-1 for the derivation.